This is Reginald again with the second question.
Suppose Bob owns 8,000 shares of Company X and 6000 shares of Company Y. The Total value of Bob's holdings of these two companies is $680,000.


Suppose Frank owns 2,000 shares of Company X and 10000 shares of Company Y. The Total value of Franks holdings of theses two companies is $680,000.

a. Write equations for Bob and Frank's Holdings. Use the variables X and Y to represent the values of shares of Company X and Y.

b. Solve for the value of a share of Company X and Company Y.

Answer by DeepaEvangeline(15)

Value of a share of Company X = X
Value of a share of Company Y = Y
Eqn for Bob's holdings is
8000X + 6000Y = 680000
Dividing the above eqn by 2000 gives
4X + 3Y = 340 --- (1)
Eqn for Frank's holdings is
2000X + 10000Y = 680000
Dividing the above eqn by 2000 gives
X + 5Y = 340 --- (2)

b) Solving for X and Y
Eqn(2) x 4 gives
4X + 20 Y = 1360 --- (3)
Eqn(3) - Eqn(1) gives
17Y = 1020 => Y = 60
Substitute the value of Y = 60 in eqn(2) gives
X + 5.60 = 340
X = 340 - 300 = 40

X = 40, Y = 60